OpenStudy (anonymous):

-6=<3(x+5)<=18 solve and graph the compound inequality, if possible

OpenStudy (anonymous):

-6 <= 3(x+5) <= 18 Distribute -6 <= 3x + 15 <= 18 Split into two equations 3x + 15 >= -6 3x + 15 <= 18 Solve for both x >= -7 x <= 1 Final inequality is -7 <= x <= 1 To graph it